# Wannier states and spin supersolid physics in the triangular antiferromagnet K2Co(SeO3)2

### v1

@@ V Quantum Spin Dynamics and Inelastic Neutron Scattering
The comparison becomes more challenging at finite fields, because the magnetic structure continuously evolves into the collinear up-up-down phase, implying that most of the spectral weight of ${\cal S}^{zz}({\bf q},\omega)$ comes from longitudinal excitations corresponding to the two-magnon continuum, while the spectral weight from transverse modes is mainly captured by the ${\cal S}^{xx}({\bf q},\omega)$ and ${\cal S}^{yy}({\bf q},\omega)$ components. While this creates a noticeable discrepancy, we can still observe in Figure 14(b) a clear energy splitting of the pseudo-Goldstone mode and the flattening of the roton minimum at $\mu_{0}H=0.25~{}\textrm{T}$. Finally, the lower band is partially recovered at $\mu_{0}H=0.5~{}\textrm{T}$ [Figure 14(c)]. To emphasize the differences with LSWT, Figure 18 presents the dynamical spin structure factor ${\cal S}^{zz}({\bf q},\omega)$ calculated within this approximation.
Figure 15: Sketch of the high-energy excitations in the Ising model at zero field. The excitation corresponds to a flip in the central blue spin changing the net magnetization in ±$\pm 1$ and the energy in (a) $J_{zz}$; (b) $2J_{zz}$; (c) $3J_{zz}$.
Lastly, we discuss the higher-energy excitations at zero field. In the pure Ising model, dispersionless single-magnon bands are predicted at energies $E=nJ_{zz}$ with $n=1,2,3$, corresponding to spin-flip excitations within a locally magnetized environment as illustrated in Figure 15. For K2Co(SeO3)2, the excitation corresponding to $n=1$ has been observed at about 3 meV in zero magnetic field, using incident neutron energy $E_{i}=7.73$ meV, as shown in Figure 16. The data, averaged over equivalent reciprocal space paths [Figure 22], reveals that this excitation becomes slightly dispersive due to the nonzero $J_{xy}$ ($\alpha\approx 0.07$), with the bottom of the branch located at the K point. The average energy of the mode is consistent with the exchange interaction $J_{zz}=3.1$ meV, as deduced from high-field magnetization data [28]. Instead of a sharp coherent mode, this excitation manifests as a gapped continuum, with its broadening attributed to strong quantum fluctuations within the Wannier sector. Similar to the lower-energy modes, a sharp excitation might exist at the lower bound of the continuum, as suggested by the shaded Gaussians in the energy-dependent cuts plotted in Figure 17.
Figure 16: (a) False-color plot of neutron scattering intensity measured on K2Co(SeO3)2 at the AMATERAS spectrometer with $E_{i}$ = 7.73 meV at $T=70$ mK and $H$ = 0, as a function of energy and wave vector transfers along high-symmetry directions in the Brillouin zone. The energy resolution is 0.24 meV at the elastic position. The data were integrated over ±$\pm 0.5$ r.l.u. in $l$ and ±$\pm 0.05$ r.l.u. in the $(h,k,0)$ plane perpendicular to the cut, and averaged over equivalent paths shown in Figure 22. No background has been subtracted. White hexagon represents the Brillouin zone boundary. $\Gamma$, K and M label high symmetry points in the reciprocal space. Labels above the top axis indicate the positions of energy-cuts shown in Figure 17.
Figure 17: Energy-cuts through the INS data shown in Figure 16 at representative wave vectors: (a) K point, (b) M point, (c) half way between $\Gamma$ and K, and (d) half way between M and $\Gamma$. The wave vectors are indicated in Figure 16 by arrows. The spectra have been averaged over equivalent wave vectors. Note the y-axis is in logarithmic scale. The inset shows the same plot with a $y$-axis on a linear scale. The shaded Gaussian represents the energy resolution. The dashed lines are empirical fits to the background.
VI Discussion
@@
In the previous sections, we demonstrated that the XXZ Hamiltonian given in Eq. (1) and the corresponding low-energy theory valid in the small $J_{xy}/J_{zz}$ limit accurately describe the low-temperature magnetic properties of K2Co(SeO3)2. While we cannot completely rule out additional interactions, the good agreement between theory and experiment suggests that such terms are likely very weak and would only become relevant at extremely low temperatures or energy scales. Therefore, at the energy scales studied, we can conclude with reasonable confidence that K2Co(SeO3)2 exhibits a spin-supersolid phase, which can be described by a nearest-neighbor XXZ Hamiltonian with strong easy-axis anisotropy. The validation of this model opens the door to exploring new questions like: What is the nature of the continuum of low-energy excitations? Can this continuum be described with a renormalized semi-classical theory by resuming higher order corrections in $1/S$ [54] or it is associated with weakly confined spinon excitations which are better described by large-$N$ approaches [55]?
@@
A few theoretical works [56, 57, 58] have recently studied a structurally similar triangular lattice antiferromagnet Na2BaCo(PO4)2, which can be described by the same easy axis XXZ Hamiltonian [13]. The ground state of Na2BaCo(PO4)2 has also been proposed to be a similar spin supersolid with the “Y” structure [13, 59], and the magnetic excitations measured by inelastic neutron scattering show a broad continuum [60]. The difference is that in Na2BaCo(PO4)2 the exchange anisotropy ratio $\alpha\approx$ 0.6 which is further away from the Ising limit compared with K2Co(SeO3)2 ($\alpha\approx$ 0.07), and the exchange interactions are much weaker with $J_{zz}$ = 0.125 meV [61, 60]. Nevertheless, we may still gain some insights on the nature of the scattering continuum in K2Co(SeO3)2 by comparing our measured spectra with the theoretically predicted ones qualitatively.
@@
One proposed scenario suggests that the spin supersolid ground state is a precursor to a Dirac spin liquid [56]. In this framework, the continuum of excitations arises from fractionalized spinons in the spin liquid state. The calculated excitation spectrum is qualitatively similar to the experimental observations: a continuum of excitations bounded by linear dispersion at the K and $\Gamma$ points. The calculated constant-energy slice of the magnetic spectrum shows excitation pockets centered not only at the $\Gamma$ and K points, but also near the M point and the wave vector halfway between $\Gamma$ and K [28]. However, at the M point, we experimentally observe a gapped, roton-like dip instead of a gapless excitation with linear dispersion. If the system is indeed in the proximity of a U(1) Dirac quantum spin liquid, one must observe a softening of the roton-mode upon approaching the “quantum melting point”.
@@
Other theoretical works offer alternative interpretations. Using tensor network methods, the magnetic excitation spectrum of the spin supersolid reveals various discrete magnon modes that are very close in energy. The observed continuum of excitations is then attributed to the finite experimental energy resolution [57, 58]. The calculated low-energy magnon modes exhibit a peculiar double magnon-roton structure: the lowest-energy magnon mode consists of a gapless Goldstone mode at the K point and a roton-like dip in dispersion at the M point, while the second-lowest energy magnon mode shows a gapped pseudo-Goldstone mode at the K point and another roton-like dip at the M point, nearly degenerate with the lowest branch. The former is attributed to fluctuations perpendicular to the plane of the ordered moment (purely transverse), while the latter is associated with fluctuations in the plane of the “Y” structure. This matches our observations in K2Co(SeO3)2 [see Figure 5 inset and Figure 11(a), (b)]. In Na2BaCo(PO4)2, the pseudo-Goldstone mode is predicted to be around 0.01 meV [58], which is extremely difficult to identify experimentally. In contrast, in K2Co(SeO3)2, thanks to the high energy resolution of AMATERAS, the pseudo-Goldstone mode as low as 0.06 meV can be resolved [Figure 11(a)].
@@
While experimentally we cannot resolve the two nearly degenerate roton-like modes predicted by theory at the M point in zero magnetic field [57, 58], their presence is clearly supported by their distinct field dependence. It has been predicted that the roton-like dip in the gapless excitation branch with the Goldstone mode vanishes quickly for increasing magnetic field, whereas that associated with the gapped pseudo-Goldstone branch persists until the field-induced quantum phase transition to the uud phase [58]. As shown in Figs. 12(b) and (c), this also agrees well with our experimental observation. The diminished visibility of the roton dip at 0.5 T may stem from its shallowing as the magnetic field increases, making it more challenging to discern experimentally. Theoretically, it has been proposed that the roton modes in the pseudo-Goldstone branch arise due to a magnon-Higgs scattering mechanism [58, 62], while those in the gapless Goldstone branch result from energy level repulsion imposed by the former. This explains its disappearance under a magnetic field, as the pseudo-Goldstone branch shifts to higher energy [see Figure 12 (c)].
@@
Finally, we discuss the gapped continuum near 3 meV. Since Na2BaCo(PO4)2 ($\alpha=0.6$) is much further from the Ising limit compared to K2Co(SeO3)2 ($\alpha=0.07$), such excitations due to spin-flips are not present in the theories mentioned above. However, in a more recent theory for K2Co(SeO3)2 [63], such an excitation is indeed reproduced. Still, the transformation of the sharp, flat, dispersive-less spin-flip excitation in a pure Ising system into a continuum as the system deviates slightly from the Ising limit arises from the quantum fluctuations within the Wannier manifold induced by $J_{xy}$. If these excitations were due to closely spaced magnon modes, the energy difference would have to fall below our experimental resolution. The low-energy excitations have also been studied within the same theoretical framework, with the roton-like dip at the M point predicted. However, the calculated spectrum is broadened by an energy resolution of 0.05 meV, making a quantitative comparison with our high-resolution data challenging.
VII Conclusion
@@
Acknowledgements.
We thank James Neuhaus for the useful discussion in the setup of SmoQyDEAC.jl. This work was Supported by a MINT grant of the Swiss National Science Foundation. Data at J-PARC were collected in Experiment no. 2023B0161. Experiments at the ISIS Pulsed Neutron and Muon Source were supported by a beamtime allocation from the Science and Technology Facilities Council under proposal no. RB2320138. L. M. Ch. was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Science and Engineering Division.
Appendix A Dynamical spin structure factor computed by LSWT
Figure 18 shows the total (left) and ${\cal S}^{zz}$ component (right) of the dynamical spin structure factor computed using linear spin-wave theory.
Figure 18: Total and ${\cal S}^{zz}$ component of the dynamical spin structure factor computed by linear spin wave theory for (a),(b) $H=0$ T, (c),(d) $\mu_{0}H=0.25$ T and (e),(f) $\mu_{0}H=0.5$ T and (g),(h) $\mu_{0}H=0.75$ T. The FWHM used in the Dirac delta functions corresponds to the experimental resolution.
Appendix B Details about QMC calculations of the DSSF
@@
$A_{\bf q}$ and $\omega^{*}$ represent fitting parameters associated with the residue and frequency of the pole, respectively. These parameters are acquired through a non-linear fitting process [see Figure 19] [64].
Figure 19: Non-linear fitting Eq. 14 for QMC data for two ${\bf q}=M$. Since we are interested in the behavior on the long imaginary time $\tau\rightarrow\beta$ a cutoff is imposed on the data fitting process.
To attack the problem of inverting Eq. 12, we employ the “Differential Evolution Algorithm” introduced and implemented in [42, 43]. After applying this procedure, the data is convoluted with the experimental resolution.
Appendix C Reciprocal paths
Figure 20: Schematic diagram showing the equivalent reciprocal space paths (a) $\Gamma$-K, (b) K-M, and (c) M-$\Gamma$, used to plot the averaged excitation spectrum in Figure 5. The shaded area is the reciprocal space covered by the detector bank at AMATERAS with $E_{i}$ = 1.69 meV. Black hexagons represent the Brillouin zone boundaries.
Figure 21: Schematic diagram showing the equivalent reciprocal space paths (a) $\Gamma$-K, (b) K-M, and (c) M-$\Gamma$, used to plot the averaged excitation spectrum in Figure 12. Shaded area is the reciprocal space covered by the detector bank at LET with $E_{i}$ = 2.15 meV. Black hexagons represent the Brillouin zone boundaries.
Figure 22: Schematic diagram showing the equivalent reciprocal space paths (a) $\Gamma$-K, (b) K-M, and (c) M-$\Gamma$, used to plot the averaged excitation spectrum in Figure 16. The shaded area is the reciprocal space covered by the detector bank at AMATERAS with $E_{i}$ = 7.73 meV. Black hexagons represent the Brillouin zone boundaries.

### v2

@@ II.6 QMC simulations
The comparison becomes more challenging at finite fields, because the magnetic structure continuously evolves into the collinear up-up-down phase, implying that most of the spectral weight of ${\cal S}^{zz}({\bf q},\omega)$ comes from longitudinal excitations corresponding to the two-magnon continuum, while the spectral weight from transverse modes is mainly captured by the ${\cal S}^{xx}({\bf q},\omega)$ and ${\cal S}^{yy}({\bf q},\omega)$ components. While this creates a noticeable discrepancy, we can still observe in Figure 18(b) a clear energy splitting of the pseudo-Goldstone mode and the flattening of the roton minimum at $\mu_{0}H=0.25~{}\textrm{T}$. Finally, the lower band is partially recovered at $\mu_{0}H=0.5~{}\textrm{T}$ [Figure 18(c)].
Figure 11(e-g) compares the QMC calculations with experimental data at the M-point, demonstrating that our simulations accurately predict the location of the lowest-energy quasiparticle peak. This agreement is further supported by the good match between the first pole and the bottom of the continuum in Figure 10.
To highlight the differences with LSWT, Supplementary Figure 6 shows the total DDSF and its ${\cal S}^{zz}$ component calculated using this approach.
III Discussion
@@
Throughout this work, we presented strong arguments supporting that the XXZ Hamiltonian (Eq. (1)) and its corresponding low-energy theory, valid in the small $J_{xy}/J_{zz}$ limit, accurately describe the low-temperature magnetic properties of K2Co(SeO3)2. While we cannot completely rule out additional interactions, the good agreement between theory and experiment suggests that such terms are likely very weak and would only become relevant at extremely low temperatures or energy scales. Therefore, at studied energy scales, we can conclude with reasonable confidence that K2Co(SeO3)2 exhibits a spin-supersolid phase, which can be described by a nearest-neighbor XXZ Hamiltonian with strong easy-axis anisotropy. We emphasize the necessity of moving beyond semiclassical approaches to accurately capture the excitation spectrum of this model. While our analysis does not explicitly determine the nature of the continuum, it clearly demonstrates that it emerges from strong quantum fluctuations.
Even more remarkable is the prospect of identifying a material that serves as an ideal platform for exploring this physics. K2Co(SeO3)2 not only enables high-resolution experimental investigations of the excitation spectrum using modern neutral spectroscopy techniques but also provides a rare opportunity to apply QMC methods, which are typically inaccessible for general frustrated magnetic systems.
Future work can now address new questions, such as: What is the nature of the low-energy excitation continuum? Can it be captured by a renormalized semiclassical theory incorporating higher-order $1/S$ corrections [55], or does it originate from weakly confined spinon excitations, better described by large-$N$ approaches [56, 57]?
@@
A few theoretical studies [58, 59, 60] have recently investigated the structurally similar triangular lattice antiferromagnet Na2BaCo(PO4)2, which is described by the same easy-axis XXZ Hamiltonian [13]. The ground state of Na2BaCo(PO4)2 has also been proposed to be a spin supersolid with a “Y” structure [13, 61], and inelastic neutron scattering measurements reveal a broad continuum of magnetic excitations [62, 59]. The key distinction lies in the exchange anisotropy ratio: Na2BaCo(PO4)2 has $\alpha\approx 0.6$, significantly further from the Ising limit than K2Co(SeO3)2 ($\alpha\approx 0.07$). Consequently, our theoretical approach is not suitable for Na2BaCo(PO4)2 due to the lack of a well-defined expansion parameter. Moreover, the exchange interactions are much weaker, with $J_{zz}=0.125\,\text{meV}$ [63, 62], which results in a lower characteristic energy scale for the magnetic exciations, making it even more challenging to resolve the structure of the excitations by neutron spectroscopy.
@@
Nevertheless, we may still gain some insights on the nature of the scattering continuum in K2Co(SeO3)2 by comparing our measured spectra with the theoretically predicted ones qualitatively. One proposed scenario suggests that the spin supersolid ground state is a precursor to a Dirac spin liquid [58]. In this framework, a Dirac spin liquid state coexists with a supersolid order. The continuum of excitations arises from fractionalized spinons in the spin liquid state. The calculated excitation spectrum is qualitatively similar to the experimental observations: a continuum of excitations bounded by linear dispersion at the K and $\Gamma$ points. The calculated constant-energy slice of the magnetic spectrum shows excitation pockets centered not only at the $\Gamma$ and K points, but also near the M point and the wave vector halfway between $\Gamma$ and K [33]. However, the excitation spectrum of the precursory Dirac spin liquid is predicted to develop a small gap due to the presence of the supersolid order. While at the M point, we experimentally observe a gapped, roton-like dip consistent with the findings in [58], the measured spectrum remains gapless at the $K$ and $\Gamma$ points.
@@
Other theoretical works offer alternative interpretations. A recent DMRG study shows that the spin supersolid phase in easy-axis triangular lattice antiferromagnets exhibit a double magnon-roton dispersion at low energy -- one branch with true Goldstone mode and the other with a gapped pseudo-Goldstone mode. Roton-like dispersion dips are present in both branches at the $M$ point and are nearly degenerate. Besides these spin-wave excitations, DMRG further predicted excitation continuum which remains even when pushing to rather high energy resolution in numeric simulations [59]. The peculiar double magnon-roton structure in the low-energy excitations is also obtained in a tensor-network renormalization study [60]. The gapless branch is attributed to fluctuations perpendicular to the plane of the ordered moment (purely transverse), while the gapped one is associated with fluctuations in the plane of the Y-structure. This agrees well with our observations in K2Co(SeO3)2 [see Figure 5 and inset]. In Na2BaCo(PO4)2, the pseudo-Goldstone mode is predicted to be around 0.01 meV [60], which is extremely difficult to identify experimentally. In contrast, in K2Co(SeO3)2, thanks to the high energy resolution of the AMATERAS spectrometer, the pseudo-Goldstone mode as low as 0.06 meV can be resolved [Figure 6(a)].
Most remarkably, with improved energy resolution, the tensor network study [60] found that the magnetic excitation spectrum of the spin supersolid displays discrete magnon modes that are very close in energy. The observed continuum of excitations [62] is then ascribed to the broadening due to finite experimental energy resolution. This stands in stark contrast to the spinon picture, where the excitation continuum emerges from fractionalized spin excitations, and points to a completely different scenario for the nature of the excitations [64].
@@
In addition, the evolution of the excitation spectrum in a magnetic field has been examined within the same framework. It has been predicted that the roton-like dip in the gapless excitation branch with the Goldstone mode vanishes quickly for increasing magnetic field, whereas that associated with the gapped pseudo-Goldstone branch persists until the field-induced quantum phase transition to the uud phase [60]. As shown in Figs. 10(b) and (c), this also agrees well with our experimental observation. The diminished visibility of the roton dip at 0.5 T may stem from its shallowing as the magnetic field increases, making it more challenging to discern experimentally. Theoretically, it has been proposed that the roton modes in the pseudo-Goldstone branch arise due to a magnon-Higgs scattering mechanism [59, 65], while those in the gapless Goldstone branch result from energy level repulsion imposed by the former. This explains its disappearance under a magnetic field, as the pseudo-Goldstone branch shifts to higher energy [see Figure 10(c)].
@@
Finally, we discuss the gapped continuum near 3 meV. Since Na2BaCo(PO4)2 ($\alpha=0.6$) is much further from the Ising limit compared to K2Co(SeO3)2 ($\alpha=0.07$), such excitations due to spin-flips are not present in the theories mentioned above. However, in a more recent theory for K2Co(SeO3)2 [31], such an excitation is indeed reproduced. Still, the transformation of the sharp, flat, dispersive-less spin-flip excitation in a pure Ising system into a continuum as the system deviates slightly from the Ising limit arises from the quantum fluctuations within the Wannier manifold induced by $J_{xy}$. If these excitations were due to closely spaced magnon modes, the energy difference would have to fall below our experimental resolution. The low-energy excitations have also been studied within the same theoretical framework, and the roton-like dip at the M point is also predicted.
@@
Methods
Sample Synthesis. Single crystals of K2Co(SeO3)2 are grown as in Ref. [35]. The crystal structure contains ABC-stacked two-dimensional triangular lattice planes of Co2+ ions separated by K layers [Figure 2]. The space group is $R$-3m, and the lattice parameters are $a=5.52$, and $c=18.52$ [66].
Inelastic neutron scattering. High-resolution inelastic neutron scattering measurements were performed using the AMATERAS cold-neutron time-of-flight spectrometer [67] in J-PARC, Japan. The incident neutron energies used are $E_{i}$ = $1.69$ and $7.73$ meV. The energy resolution at elastic scattering is $23$ and $240$ $\mu$eV, respectively. The data were collected by rotating the sample by $160$ degrees with $1$ degree per step, and analyzed using the Horace package [68]. Inelastic neutron scattering measurements in a magnetic field were carried out using the LET cold-neutron time-of-flight spectrometer in ISIS facility, Rutherford Appleton Laboratory, United Kingdom. The incident neutron energy is fixed at $E_{i}$ = 2.15 meV. The spectrometer is operated in the high-flux mode, with chopper frequency 240/120 Hz, yielding an energy resolution of 40 $\mu$eV at elastic scattering. The data were obtained by rotating the sample by 90 degrees with 1 deg. per step. A single crystal sample of 425 mg and 1.23 grams were used for AMATERAS and LET, respectively. They are mounted on a copper holder and cooled using a 3He/4He dilution refrigerator. At LET, a magnetic field is applied along the [0,0,1] direction using a vertical-field cryomagnet. In both cases the sample was oriented with [1,0,0] and [0,1,0] directions in the horizontal plane.
Specific heat measurements. The specific heat measurements were performed using a relaxation method by a Quantum Design Physical Property Measurement System on a 0.05 mg single-crystal sample for 0.12--4 K with a dilution fridge insert and on a 3.95 mg one for 2--300 K.
LSWT and QMC calculations. All LSWT calculations were performed using Sunny.jl (v0.7) [69]. Worldline QMC calculations were performed using a modified directed-loop algorithm [70, 71] for system sizes of $36\times 36$ (thermodynamic properties) and $72\times 72$ (spin dynamics) with $N_{\tau}=200$ imaginary-time-steps. To verify the correctness of the QMC code, exact diagonalization calculations were performed on small clusters.
@@ Methods
$A_{\bf q}$ and $\omega^{*}$ represent fitting parameters associated with the residue and frequency of the pole, respectively. These parameters are acquired through a non-linear fitting process [72].
To attack the problem of inverting Eq. 11, we employ the “Differential Evolution Algorithm” introduced and implemented in SmoQyDEAC.jl (v1.1.4) library [73, 74]. After applying this procedure, the data is convoluted with the experimental resolution.
Data availability
All data are available upon reasonable request to the corresponding author. The data collected at LET, STFC ISIS Neutron and Muon Source, is available at https://doi.org/10.5286/ISIS.E.RB2320138.
Code availability
Sunny.jl (v0.7) and SmoQyDEAC.jl (v1.1.4) are publicly available on their respective GitHub repositories. The QMC code used for this study may be made available to qualified researchers upon reasonable request from the corresponding author.
Acknowledgements.
We are grateful to M. E. Zhitomirsky for directing us to relevant prior literature and to A. Läuchli for illuminating discussions. We have incorporated new references to appropriately acknowledge these contributions. The authors are grateful to Hao Zhang for advice on using Sunny.jl. We thank James Neuhaus for the useful discussion in the setup of SmoQyDEAC.jl. This work was Supported by a MINT grant of the Swiss National Science Foundation. Data at J-PARC were collected in Experiment no. 2023B0161. Experiments at the ISIS Pulsed Neutron and Muon Source were supported by a beamtime allocation from the Science and Technology Facilities Council under proposal no. RB2320138. L. M. C. was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Science and Engineering Division.
Author contributions
The samples used in this study were synthesized by M.Z. and Z.Y. Thermodynamic measurements were performed by M.Z. and S.G.. Inelastic neutron scattering experiments were carried out by M.Z., V.R. and A.Z. with assistance from N.M., S.O. and C.B. QMC calculations were performed by L.M.C and C.D.B. Y.K. developed and tested the QMC code. The manuscript was written by M.Z., L.M.C., A.Z. and C.D.B. with input from all authors.
Competing interests
The authors declare no competing interests.